-6x+6y=6 -6x+3y=-12
Answer by TimothyLamb(4379)

---
-6x+6y=6
-6x+3y=-12
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution by matrix method:
x = 5
y = 6
---
now by elimination method:
---
-6x+6y=6
(-6x+3y=-12) * -1
---
-6x+6y=6
6x-3y=12
---
add the system:
---
3y = 18
y = 6
---
-6x + 6(6) = 6
-6x + 36 = 6
-6x = -30
x = 5
---
solution by elimination method:
x = 5
y = 6
---
both methods agree !
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php